1. what is 2/3=m/42
2. 6.3=0.9y y=
3. 6= z/5.9 z=

Answer by Alan3354(69443)

1. what is 2/3=m/42
It's an equation in m.
------------------
2. 6.3=0.9y
y = 6.3/0.9
You can divide that easily and get y = 7
Sometimes it's better to eliminate fractions first.
6.3=0.9y
Multiply by 10
63 = 9y
y = 63/9
y = 7
--------------
3. 6= z/5.9
Multiply by 5.9
z = 6*5.9
z = 35.4